OpenStudy (anonymous):

f(x)=3x-1 and g(x)=x^2+4. Find f(g(6))

OpenStudy (anonymous):

Plug in 6 into g(6) and then plug in the answer of g(6) into f(x)

OpenStudy (unklerhaukus):

\[f(g(x))=f(x^2+4)=3(x^2+4)-1=\]

OpenStudy (anonymous):

I keeep getting the answer wrong.

OpenStudy (unklerhaukus):

what are you getting?

OpenStudy (anonymous):

g(6) = (6)^2 + 4 = 36 + 4 = 40. f(40) = 3(40) - 1 = 120 - 1 = 119

OpenStudy (anonymous):

how do i type that answer

OpenStudy (anonymous):

I am doing practice questions on the computer and no matter how i write it it says its wrong?@

OpenStudy (unklerhaukus):

\[f(g(6))=f(6^2+4)=3(6^2+4)-1=\]

OpenStudy (raden):

f(g(6)) = (fog)(6) @UnkleRhaukus and @vintageballad are right

OpenStudy (anonymous):

i didnt know if there was another way to write the answer

OpenStudy (unklerhaukus):

did you get a single value when simplified ?

OpenStudy (raden):

now, u knowed that :)

OpenStudy (anonymous):

(fog)?

OpenStudy (anonymous):

yes :0)

OpenStudy (unklerhaukus):

\[(f\circ g)(x)=f(g(x))\]

OpenStudy (anonymous):

I think I am confusing myself

OpenStudy (unklerhaukus):

look at vintageballad's work